OpenStudy (anonymous):

Use natural logarithms to solve the equation. Round to the nearest thousandth. 6e4x - 2 = 3 A. –0.448 B. 0.327 C. 0.067 D. –0.046

OpenStudy (anonymous):

what is the e thing, also show work so i can understand

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@JakeV8

jimthompson5910 (jim_thompson5910):

\[\Large 6e^{4x}-2=3\] ???

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

Start off by adding 2 to both sides

OpenStudy (anonymous):

\[6e ^{4x}=5\]

jimthompson5910 (jim_thompson5910):

now divide both sides by 6

OpenStudy (anonymous):

\[e ^{4x}= .8333333333\]

jimthompson5910 (jim_thompson5910):

then convert to logarithmic form

jimthompson5910 (jim_thompson5910):

this gives you 4x = ln(0.833333333)

jimthompson5910 (jim_thompson5910):

the next thing to do is evaluate the right side finally, divide both sides by 4 and round

OpenStudy (anonymous):

no im confused i dont know how to put this in my calculator

jimthompson5910 (jim_thompson5910):

there is an "ln" button (LN)

OpenStudy (anonymous):

nope

jimthompson5910 (jim_thompson5910):

probably near the Log button

jimthompson5910 (jim_thompson5910):

what kind of calculator do you have

OpenStudy (anonymous):

TI-84

jimthompson5910 (jim_thompson5910):

under the log button is the LN button

jimthompson5910 (jim_thompson5910):

type LN, then enter 0.8333333 and hit enter

OpenStudy (anonymous):

not on mine its sto

jimthompson5910 (jim_thompson5910):

weird, guess I'm looking at the wrong version?

OpenStudy (anonymous):

o i found it

OpenStudy (anonymous):

-.1823215969

jimthompson5910 (jim_thompson5910):

now divide that by 4

OpenStudy (anonymous):

-.0455803992

jimthompson5910 (jim_thompson5910):

then round to 3 decimal places

OpenStudy (anonymous):

-0.046??

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

Another way to do it \[\Large 6e^{4x}-2=3\] \[\Large 6e^{4x}=3+2\] \[\Large 6e^{4x}=5\] \[\Large e^{4x}=\frac{5}{6}\] \[\Large 4x=\ln\left(\frac{5}{6}\right)\] \[\Large 4x\approx-0.18232155679\] \[\Large x\approx\frac{-0.18232155679}{4}\] \[\Large x\approx-0.045580389\] \[\Large x\approx-0.046\]

OpenStudy (anonymous):

thanks you for your help

jimthompson5910 (jim_thompson5910):

oh wait, nvm, I'm thinking of something else...but what I showed just now is what you just did

OpenStudy (anonymous):

o so im wrong

jimthompson5910 (jim_thompson5910):

no you are 100% correct

OpenStudy (anonymous):

o ok lol

jimthompson5910 (jim_thompson5910):

I just thought my steps were different from yours (when they aren't really)